OpenStudy (anonymous):

Find the values of x and y that maximize the objective function P =3x +2y for the graph. What is the maximum value?

OpenStudy (anonymous):

x = 9, y = 0 is the maximum value for this function. hence it is 27.

OpenStudy (anonymous):

Sure why not :)

OpenStudy (anonymous):

If you simplify the second equation, you get -3(3x+3y) = -30 which is 3x+3y = 10. Both the equations are the exact same values. hence any value working for the first would work for the second.

OpenStudy (anonymous):

So there is no specific answer?

OpenStudy (anonymous):

@Mathbreaker

OpenStudy (anonymous):

No not really. you cant conclude from these equations cause they're the same. however they basically also mean that x + y = 10/3 now we know that 1/3 + 9/3 = 10/3 hence this could be x and y but also 2/3 + 8/3 = 10/3 so this could work as well. even 1.3/3 + 8.7/3 = 10/3... This concludes infinite solutions within the system.

OpenStudy (anonymous):

Thank You! Last One?

OpenStudy (anonymous):

@Mathbreaker

OpenStudy (anonymous):

x = 3.4, y = 0.4, z = -4

OpenStudy (anonymous):

Thanks!! If you can, can you quickly tell me how you got that....?

OpenStudy (anonymous):

z= 7 - 3x - 2y replace it in the other 2 equations you get 2 simultaneous equations for x and y solve them then when you have x and y, solve one equation for z. I wouldn't written more on this but i really need to go. I hope you get it from this...:)

OpenStudy (anonymous):

Thank that really helped!

OpenStudy (anonymous):

You're welcome :)